OpenStudy (anonymous):

cosh2x = cosh^2x + sinh^2x Prove this identity!

OpenStudy (anonymous):

\[\cosh2x = \cosh^2x + \sinh^2x\]

OpenStudy (solomonzelman):

cosh^2x, do you mean cos (h^2x) ?

OpenStudy (anonymous):

no this is hyperbolic functions lol

OpenStudy (anonymous):

Calc 2, wio knows what's up

OpenStudy (solomonzelman):

Dang, ....

OpenStudy (anonymous):

Okay, so they probably don't want you to use Osborn's rule here.

OpenStudy (anonymous):

nope lol

OpenStudy (solomonzelman):

wio - calculator.

OpenStudy (anonymous):

I know that \[\cosh^2x = (\frac{ e^x+e^-x }{ 2 })^2\] and \[\sinh^2x = (\frac{ e^x-e^-x }{ 2 })^2\]

OpenStudy (anonymous):

-x should be ^

OpenStudy (anonymous):

So I think we should make cosh^2x on the left side basically = \[(e^x+e^-x/2)^2 \]

OpenStudy (anonymous):

What ya think wio?

OpenStudy (anonymous):

It's simple. Expand out the left hand side. Expand out the right hand side.

OpenStudy (anonymous):

All the expanding you do on the right hand side can be done in reverse, as if you were really good at factoring, for the actual proof.

OpenStudy (anonymous):

i always forget cosh^2x = (coshx)(coshx)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

When you expand \(\cosh^2(x)\) notice:\[ \cosh^2(x) = \frac{e^{2x}+2e^0+e^{-2x}}{4} \]

OpenStudy (anonymous):

yup yup

OpenStudy (anonymous):

Just wrote that down as well

OpenStudy (anonymous):

So when doing things in reverse, you have to subtract 2 and add 2 at the same time.

OpenStudy (solomonzelman):

wio is the best! I told you! Dr calculator....!

OpenStudy (anonymous):

@wio is e^x * -e^-x = -1?

OpenStudy (solomonzelman):

e=2.7

OpenStudy (solomonzelman):

right

OpenStudy (anonymous):

Yeah, that is correct.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

What exactly does cosh2x equal to?

OpenStudy (anonymous):

is it just e^2x+e^-2x/2 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

mhm ok, let me see if I got it now

OpenStudy (anonymous):

nope..

OpenStudy (anonymous):

@wio I did something wrong, I'm getting \[e^{2x} + e^{-2x} + 2\]

OpenStudy (anonymous):

You can rewrite the right side as \(2sinh^2x +1\) and save yourself a lot of stupid multiplying

OpenStudy (anonymous):

I really wish you mentioned that before lmao

OpenStudy (anonymous):

I'll give it a shot and see what I get this time >.<

OpenStudy (anonymous):

Well what @SACAPUNTAS said is true, but they might not want that depending on the rigor involved.

OpenStudy (anonymous):

This is why I hate identities.

OpenStudy (anonymous):

\[\begin{split} \cosh^2(x) + \sinh^2(x) &=\left(\frac{e^{x}+e^{-x}}{2}\right)^2+\left(\frac{e^{x}-e^{-x}}{2}\right)^2\\ &=\frac{e^{2x}+2+e^{-2x}}{4}+\frac{e^{2x}-2+e^{-2x}}{4}\\ &=\frac{e^{2x}+e^{2x}+\cancel{2-2}+e^{-2x}+e^{-2x}}{4}\\ &=\frac{2e^{2x}+2e^{-2x}}{4}\\ &=\frac{e^{2x}+e^{-2x}}{2}\\ &=\cosh(2x) \end{split} \]

OpenStudy (anonymous):

Write this backwards.

OpenStudy (anonymous):

Ah very very nice, thank you so much Wio!!!

OpenStudy (anonymous):

I found my mistake as well haha, I added the 2s >_>